5x + 7y + 8z = 20
8x + 3y + 10z = 70
9x + 4y + 9z = 80
I need system solution for the above system of equations by using Gaussian Elimination Method. Thanking you in anticipation.

Answer by Edwin McCravy(20056)

The solution to that system of equations is as 
follows using technology:

x = 1250/101

y = -410/101

z = -170/101

Since I do not believe that any teacher would give
a problem with such a terrible solution to do by
hand, I believe that you have made an error in
copying the system of equations.  Please check.

Edwin